OpenStudy (anonymous):

Y=2^x , Dervative(4) , i get 32 is it correct?

OpenStudy (jtvatsim):

Derivative of exponential is given by Y = 2^x -> Y' = ln 2 * 2^x Thus, Y'(4) = ln 2 * 2^4 = 16 ln 2. A bit of a nasty number.

OpenStudy (jtvatsim):

If you need the general equation it is \[\frac{d}{dx} (a^x ) = (\ln a) \cdot a^x\]